OpenStudy (fly_g):

Use the function to answer the question. f(x) = 4|x + 1| – 6 What is the vertex of the graph of f(x)? (–1, –6) (–1, 6) (1, –6) <--what i picked (4, –6)

OpenStudy (solomonzelman):

So for parent function, \(\large\color{black}{ \rm y=4\left| x \right| }\) the vertex is (0,0). Your function though is shifted. Find out the shift based on the rules bloew, for me please. Rules of \(\large\color{black}{ \rm shifts }\) from \(\large\color{black}{ \rm f(x) }\) to \(\large\color{black}{ \rm g(x) }\). \(\large\color{black}{ \rm f(x)=4\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=4\left| x \color{blue}{ -~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm {c} }\) units to the \(\normalsize\color{blue}{ \rm right }\). \(\large\color{black}{ \rm f(x)=4\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=4\left| x \color{blue}{ +~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm {c} }\) units to the \(\normalsize\color{blue}{ \rm left }\). \(\large\color{black}{ \rm f(x)=4\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=4\left| x \right| \color{blue}{ +~\rm{c} }}\) \(\large\color{blue}{ ~\rm {c} }\) units \(\normalsize\color{blue}{ \rm up }\). \(\large\color{black}{ \rm f(x)=4\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=4\left| x \right| \color{blue}{ -~\rm{c} }}\) \(\large\color{blue}{ ~\rm{c} }\) units \(\normalsize\color{blue}{ \rm down }\). if you have questions please ask.